OpenStudy (anonymous):

A grocer wants to make a 10-pound mixture of peanuts and cashews that he can sell for $4.75 per pound. If peanuts cost $4.00 per pound and cashews cost $6.50 per pound, how many pounds of cashews should he use? 3 4 6 7

OpenStudy (anonymous):

do u know how to do this

OpenStudy (jackellyn):

Let C and P be the pounds of each in the mix 6.50 C + 4.00P = 47.50 (dollars) C + P = 10 (pounds) You do the solving. Try substituting 100 - C for P in the first equation, and then solve for C. Use back substitution.

OpenStudy (anonymous):

@johhjohn Can you do it?

OpenStudy (anonymous):

im tryin

OpenStudy (anonymous):

More hints :) 6.5 c + 4 ( 10 - c) = 47.5

OpenStudy (anonymous):

Does it help?

OpenStudy (jackellyn):

my bad, 10-c not 100-c, type-o

OpenStudy (anonymous):

more 2.5c=7.5 =c=3

OpenStudy (anonymous):

its cool an thats the part that rly messed me up

OpenStudy (anonymous):

Yep, 3 lbs of cashew

OpenStudy (anonymous):

thax

OpenStudy (anonymous):

Can you find the pounds of peanut ?

OpenStudy (anonymous):

Hint: Total mixed pounds is 10 lbs :)

OpenStudy (anonymous):

Even it's not in the question but you should know for your own good!

OpenStudy (anonymous):

ya 7 right

OpenStudy (anonymous):

The essential concept here is to construct the equation: 6.5 c + 4 ( 10 - c) = 47.5

OpenStudy (anonymous):

wat u mean

OpenStudy (anonymous):

The calculation part just a trivial step after you write down the equation!

OpenStudy (anonymous):

ook